# Potential-Barrier Affinity Effect in Solid Systems

### 17 Mar 2026 v3

@@ Methods
Figure 6: The electron density distribution functions for near-free electrons ($\rho_{F}$) and bound-state electrons ($\rho_{B}$), where $\rho_{F}$ and $\rho_{B}$ are calculated according to the unbound-state bands ($V_{M}<\epsilon_{nk}<E_{F}$) and bound-state bands ($\epsilon_{nk}<V_{M}$). The isosurfaces of 0.008, 0.005, 0.008, 0.01, 0.04, 0.1, and 0.05 Å-3 are used for Ca2N, Ba2N, Sr2N, Li5Si, Li6Al, Li8Au, and Li4Rh, respectively.
@@
To further establish that IAEs arise from near-free-electron (NFE) behavior, we calculated the partial electron density from band decomposition for the electrides in Fig. 2(d) of the maintext, as well as the representative mayenite [Ca24Al28O64][11]. he results, presented in the Figure 6 and Figure 7, reveal that the NFE density accumulates predominantly in the interstitial region, in agreement with the findings for Na-hP4.
@@
Although early theoretical models attributed the interstitial-potential well to Pauli exclusion from core electrons[21, 2], no accurate or quantitative theory exists within the fermionic quasiparticle framework to describe the Pauli potential. Here, to comprehensively analyze the impact of Pauli exclusion in the PBA effect, we mapped the noninteracting fermions (Kohn-Sham system) to the reference noninteracting bosons, where the bosons have the same charge, mass, point-like shape, chemical potential, and density as the electrons[39, 40], as shown in Fig. 9(a). In KSDFT, the electronic total energy of the interacting systems is defined as:
@@
where the chemical potential is set as the Fermi energy level of KSDFT in this work ($\mu=E_{F}$). To include the interaction between nuclear and valence electrons, we performed the all-electron (AE) KSDFT calculations using the full-potential linearised augmented-planewave method[43] by the ELK-v10.6.2 code[44]. The AE effective potentials and electron densities are presented in Figure 9(c-d) for Na-hP4 $\bar{1}1$], Al-FCC [111], and Diamond [111] directions. In this framework, the collective quantum state exhibits bound-state behavior under the all-electron effective potential containing Pauli interaction. The results completely overwhelm the behavior of fermionic quasiparticles, making it exceptionally difficult to analyze electron behavior. For the AE KSDFT calculations, we applied $18\times 18\times 11$ and $18\times 18\times 18$ $k$-grid meshes for the ground-state calculations of Na-hP4 and Al-FCC/Diamond, respectively. The potential and electron density are represented in the dense real-space grids of $120\times 120\times 120$. The eight-order finite-difference method[45] is adopted for evaluating vW kinetic potential.

### 14 Apr 2026 v4

@@ Methods
Figure 6: The electron density distribution functions for near-free electrons ($\rho_{F}$) and bound-state electrons ($\rho_{B}$), where $\rho_{F}$ and $\rho_{B}$ are calculated according to the unbound-state bands ($V_{M}<\epsilon_{nk}<E_{F}$) and bound-state bands ($\epsilon_{nk}<V_{M}$), respectively. The isosurfaces of 0.008, 0.005, 0.008, 0.01, 0.04, 0.1, and 0.05 Å-3 are used for Ca2N, Ba2N, Sr2N, Li5Si, Li6Al, Li8Au, and Li4Rh, respectively.
@@
To further establish that IAEs arise from near-free-electron (NFE) behavior, we calculated the partial electron density from band decomposition for the electrides in Fig. 2(d) of the maintext, as well as the representative mayenite [Ca24Al28O64][11]. The results, presented in the Figure 6 and Figure 7, reveal that the NFE density accumulates predominantly in the interstitial region, in agreement with the findings for Na-hP4.
@@
Although early theoretical models attributed the interstitial-potential well to Pauli exclusion from core electrons[21, 2], no accurate or quantitative theory exists within the fermionic quasiparticle framework to describe the Pauli potential. Here, to comprehensively analyze the impact of Pauli exclusion, we mapped the noninteracting fermions (Kohn-Sham system) to the reference noninteracting bosons, where the bosons have the same charge, mass, point-like shape, chemical potential, and density as the electrons[39, 40], as shown in Fig. 9(a). In KSDFT, the electronic total energy of the interacting systems is defined as:
@@
where the chemical potential is set as the Fermi energy level of KSDFT in this work ($\mu=E_{F}$). To include the interaction between nuclear and valence electrons, we performed the all-electron (AE) KSDFT calculations using the full-potential linearised augmented-planewave method[43] by the ELK-v10.6.2 code[44]. The AE effective potentials and electron densities are presented in Figure 9(c-d) for Na-hP4 $\bar{1}1$], Al-FCC [111], and Diamond [111] directions. In this framework, the collective quantum state exhibits bound-state behavior under the all-electron effective potential containing Pauli interaction. The results completely overwhelm the behavior of fermionic quasiparticles, making it exceptionally difficult to analyze electron behavior. Therefore, the PBA effect derived from our approach offers a mathematically rigorous alternative that captures the essential physics. For the AE KSDFT calculations, we applied $18\times 18\times 11$ and $18\times 18\times 18$ $k$-grid meshes for the ground-state calculations of Na-hP4 and Al-FCC/Diamond, respectively. The potential and electron density are represented in the dense real-space grids of $120\times 120\times 120$. The eight-order finite-difference method[45] is adopted for evaluating vW kinetic potential.